# Ten Years of ZMap

### v1

@@ 4.2. Sharding
In 2014, Adrian et al. (adrian2014zippier, ) introduced a mutex-free sharding mechanism that enabled spreading address generation across multiple senders without synchronization. This enabled a single Internet-wide scan to be split across multiple machines and improved performance by allowing multiple send threads on one machine to operate independently. As Mazel et al. noted when they showed that ZMap can be fingerprinted through its IP generation method (mazel2019identifying, ), we further shifted the sharding approach in 2017, as described below.
@@
Interleaved Sharding. Sharding was originally implemented with each shard iterating by $N$ steps at a time with offset starting elements. For example, for $N$ shards, each shard $n\in[1\mathrel{{.}\,{.}}N]$ iterates by multiplying the current element by $g^{N}$, but begins iteration at $g^{n}$ (Figure 6(a)). With multiple threads in place, each shard $n$ is then further split into $T$ subshards, with each subshard iterating by $g^{NT}$ offset by $g^{n+tN}$, where $n$ is the shard index and $t$ is the thread index.
While conceptually simple, the approach requires calculating the end point of each shard to know when to stop iterating. $NT$ is not guaranteed to cleanly divide $p-1$, and so a shard might not repeat its initial value until it finishes iterating the full group (rather than just the shard). Unfortunately, the last index of a shard does not have a closed form expression, and we found that calculating it is prone to off-by-one errors. After repeated such correctness issues, we switched to a simpler sharding mechanism.
@@ 4.3. Packet Construction
TCP options affect packet size and therefore scan rate. However, including the MSS option alone finds the vast majority of services (over 99.99% of services on TCP/80) and remains under the minimum size of an Ethernet frame, continuing to support the maximum 1.488 Mpps line rate of a 1 GbE link. Using the Windows or Linux packet layouts finds very slightly more services but reduces send rates to 1.389 and 1.276 Mpps, respectively. While filtering packets based on TCP options could be due to defensive mechanisms attempting to block scanning, removing the easily identified static IP ID value of ZMap probes appears to have no impact on scan hit rates. We performed three scans of 10% of IPv4 on TCP/80 in April 2024 with a static IP ID and with a random per-packet IP ID and find that the difference in hit-rate between the random and static IP IDs is not statistically significant. In early 2024, ZMap changed to using random per-probe IP IDs.
Figure 7. Hitrates for Varying TCP Options—TCP SYN probes without any TCP options, as originally sent by ZMap, find 1.5–2.0% fewer services on TCP/80 than including any options. Mimicking common OSes maximizes coverage.
@@ 5. Lessons and Recommendations
Based on our experiences maintaining ZMap and feedback we received from other users, we offer the following lessons and recommendations for future measurement tools.
@@
Nearly 50 years later, we cannot agree more with this guidance. It is impossible to fully predict how researchers will use measurement tools or the environments where they will operate (e.g., running ZMap on 16-bit embedded systems). ZMap was originally envisioned as a framework where researchers would build customized Scan Modules and Output Modules for service follow up. In practice, the vast majority of researchers use ZMap for service discovery and pipe results to secondary tools for investigation or storage. The output modules ZMap included for connecting to specific databases (e.g., Redis) became liabilities, requiring upkeep and complicating testing and packaging. In time, we removed them, opting to support only Text, CSV, and JSON Lines output.
Recommendation: Build small, simple, easy-to-understand, easy-to-use, and easy-to-test measurement tools that can be creatively assembled, rather than complex applications or software frameworks. Build applications that do one thing well. Continuously output results on a per-record/per-line basis when possible. Avoid proprietary formats and standardize output on well-worn interfaces like JSON Lines (jsonlines, ) or CSV, even if those formats are less computationally efficient; interoperability with other tools is more important.
Usability. ZMap was not the first fast, asynchronous network scanner: Unicornscan (unicornscan, ), Scanrand (scanrand, ), and IRLscanner (leonard2010demystifying, ) were released years prior (though Unicornscan and Scanrand were both unknown to the originating ZMap team at the time). ZMap was more successful than these tools due to its greater ease of use: it enabled researchers to scan the Internet from a single machine by running a single command.
Recommendation: Obsess over ease of installation, usage, and troubleshooting. It is better to have a tool that is easy to use and less full featured than vice versa.
Library and Command Line Wrapper. It is natural to build an application where the command-line interface, application configuration, and operation are intermingled, since most tools are first used via the command line and this is the path of least resistance. However, this design will limit a tool’s potential to be integrated into larger systems. Automated or continous measurements, such as the scans that power services like Censys (durumeric2015search, ), are cumbersome to control from the command line and more suited to a library interface.
@@
Data, Metadata, and Logs. Unix has two standard output streams: Standard Output (stdout) and Standard Error (stderr). We argue that, in contrast, measurement tools should produce three types of output: (1) Data [Standard Output], (2) Logs [Standard Error], and (3) Metadata. Given the amount of data produced, it is difficult to tell whether an experiment is operating as expected without analyzing metadata in real time. In addition, tracking as much information about the execution (e.g., time, version of software, configuration parameters, and environment) will help to later interpret, troubleshoot, or reproduce results. Logs are helpful for human debugging, but they are not inherently designed to be machine-parsable, which is needed for monitoring long-running experiments. Ultimately, we extended ZMap to produce four output streams: (1) data, (2) logs, (3) real-time updates (e.g., packets sent, received, dropped per second), and (4) machine-readable metadata at completion.
@@
Recommendation: Design measurement tools to produce separate streams of data, metadata, and debug logs. Do not mix these streams, since this complicates downstream processing. Be liberal in what environment and execution information is included in scan metadata, as it is difficult to know a priori what will be useful. Adopt a logging library that supports multiple log levels, and use debug-level logging liberally to enable future troubleshooting.
@@
Static Types and Output Schema. JSON and CSV provide considerable flexibility when encoding data. For example, in JSON, objects can have dynamic keys, and values are not typed. However, downstream applications and databases usually do not support this flexibility, and it is easy to create valid JSON records that are painful to process.
@@
Recommendation: Define a schema for the data you output. Ensure that each field uses a single, well defined type and that the type of one field does not depend on the value of another field. Avoid maps with dynamic keys, and instead use lists of a static document type. Consider using a tool like JSON Schema (json-schema, ) or ZSchema (zschema, ) to document the structure of output data.
@@
Versioning and Releases. We released ZMap far too infrequently, repeatedly falling victim to wanting to include one more feature or bugfix in each release. In particular, ZMap 3.0 was released nearly eight years after the previous release of ZMap 2.1.1. Unfortunately, this meant that most users were either using long out-of-date releases or running HEAD. This made debugging difficult and prevented users from describing which ZMap version they used.
@@
Recommendation: Follow the Semantic Versioning Specification (semver, ) religiously. Focus on making regular versioned releases rather than trying to finish a preset amount of work.
@@
Language Choice. When we wrote ZMap, C and C++ were the only practical high-performance systems programming languages available. It is deceivingly easy to implement software in C and to convince oneself that it is possible to safely write C code. Empirical evidence overwhelmingly says the opposite (gaynor2020science, ), and even the White House now recommends moving towards memory safe languages (whitehouse2024memory, ). Network parsers are particularly hard to implement safely and must be resistant to attacker-controlled and malformed data (chromeruleoftwo, ). Over the years, ZMap has had multiple regressions that caused incorrect or incomplete measurements, as well as multiple memory safety bugs, all of which could have been avoided with a different choice of languages. Our team has also found that memory safety concerns make it harder to review external contributions for safety and correctness, which has reduced the rate at which we can merge improvements.
@@
Recommendation: Responsibly develop measurement tools in modern, memory-safe languages like Go and Rust. While Rust has a relatively steep learning curve, its safety and runtime performance make it ideal for performance-critical applications. Go’s performance, while less than Rust’s, is sufficient for many measurement applications (e.g., ZGrab (zgrab2, ) and ZDNS (zdns, )), and the language’s simple syntax and parallelism-oriented architecture make it particularly suited for quickly developing network measurement tools.
6. Conclusion
The most exciting aspect of building ZMap has been watching how other researchers have used it in a myriad of unpredicted but valuable ways. We are sincerely thankful to everyone who has contributed, pushing the tool close to being feature complete. While the development of new ZMap features has slowed, we are excited to continue to expand the ecosystem of tools that work with ZMap (e.g., ZDNS (zdns, ) and ZGrab (zgrab2, )) and thereby to enable an even broader range of measurement uses. Many of the lessons we learned from maintaining ZMap may seem obvious today, but we did not have the foresight originally. We hope that our retrospective analysis will help the community build an even richer and more reliable ecosystem of Internet measurement tools moving forward.
Appendix A Ethics
Our work is primarily a metareview of prior studies and a presentation of lessons learned from ZMap. When conducting our own experiments to validate changes made to ZMap, we followed the original guidelines set forth by Durumeric et al. in 2013 (durumeric2013zmap, ). However, revisiting the release of ZMap, this work also provides us with an opportunity to revisit the original recommendations for how best to conduct active Internet-wide scans (durumeric2013zmap, ):
@@ Appendix A Ethics
    Coordinate closely with local network admins to reduce risks and handle inquiries.
  2. (2)
    Verify that scans will not overwhelm the local network or upstream provider.
  3. (3)
    Signal the benign nature of the scans in web pages and DNS entries of the source addresses.
  4. (4)
    Clearly explain the purpose and scope of the scans in all communications.
  5. (5)
    Provide a simple means of opting out, and honor requests promptly.
  6. (6)
    Conduct scans no larger or more frequent than is necessary for research objectives.
  7. (7)
    Spread scan traffic over time or source addresses when feasible.
Overall, we believe that these recommendations remain a sound set of considerations when conducting active Internet measurement. However, we also encourage the Internet measurement community to treat these recommendations as a set of good practices for most research, not as a set of requirements nor as the basis for having conducted research ethically. For example, there may be some situations where scans may be better performed unattributed, or where opting out certain networks could invalidate an operation (e.g., when tracking a specific threat actor). Institutions have adopted different practices for validating opt-out requests . Given that IP addresses ownership can change, it may make sense to expire opt-out requests.
Beyond these original recommendations, we additionally recommend that researchers consider the following:
  1. (1)
    Investigate whether existing datasets (e.g., Censys Research Access and Rapid7 Open Data) can suffice. In many cases, existing datasets can provide higher coverage of services across a large number of ports than researcher-operated scans while simultaneously reducing the aggregate bandwidth used by scans.
@@
    Validate how handshakes will appear in the logs of scanned systems. For example, some SSH handshakes that do not attempt to authenticate still show up as failed authentication attempts causing undue worry for operators monitoring their logs.
We offer the following, updated set of best practices:
@@
    Minimize Internet Impact. While Internet scanning is a powerful methodology, it can also affect systems and create work for operators. Consider whether existing open source datasets provide the data you need. If you do perform scans, conduct scans no larger or more frequent than is necessary for research objectives. Publish data for future researchers.
@@
    Signal Intent. When possible, publish reverse DNS entries, IP WHOIS records, and an informational website that describes the benign nature of scans and how to contact the responsible researchers.
@@
    Provide An Opt-Out Mechanism. Provide a simple mechanism for operators to request to be opted out of future scans. Indicate the IP ranges you use for scanning so that operators can drop research traffic.
@@
    Proactively Investigate Effects. Run any newly developed scanning code against your own systems to ensure that you understand how scans might affect devices and appear in system logs. Start with small experiments before completing full Internet scans in case your scanner causes unexpected problems.
@@
    Coordinate Locally. Coordinate with your local IT and security teams to reduce the risk of overwhelming local networks, as well as to ensure that they know how to handle any inbound inquiries.
@@
    Disclose Results. When conducting security studies, consider how you can improve the security of the systems you have scanned. Responsibly disclose security problems you find and consider notifying individual system owners who are vulnerable.

### v2

@@ 4.2. Scan Sharding
In 2014, Adrian et al. (adrian2014zippier, ) introduced a mutex-free sharding mechanism for ZMap’s address generation. This enabled scans to be split across machines and improved performance by allowing multiple send threads on one machine to operate independently. As Mazel et al. noted when they showed that ZMap can be fingerprinted through its IP generation method (mazel2019identifying, ), we shifted the sharding approach in 2017. Given that other work is analyzing ZMap’s address generation, we describe the two approaches:
@@
Interleaved Sharding. Sharding was originally implemented with each shard iterating by $N$ steps at a time, offset by one step. For $N$ shards, each shard $n\in[1\mathrel{{.}\,{.}}N]$ iterates by multiplying the current element by $g^{N}$, but begins iteration at $g^{n}$ (Figure 6(a)). With multiple threads in place, each shard $n$ is then further split into $T$ subshards, with each subshard iterating by $g^{NT}$ offset by $g^{n+tN}$, where $n$ is the shard index and $t$ is the thread index. While conceptually simple, the approach requires calculating the end point of each shard to know when to stop iterating. $NT$ is not guaranteed to cleanly divide $p-1$, and so a shard might not repeat its initial value. Unfortunately, the last index of a shard does not have a closed form expression and we found that calculating it is prone to off-by-one errors. After repeated correctness issues, we switched to a simpler mechanism.
@@ 4.3. Packet Construction
TCP options affect packet size and therefore scan rate. However, including the MSS option alone finds the vast majority of services (over 99.99% of services on TCP/80) and remains under the minimum size of an Ethernet frame, continuing to support the maximum 1.488 Mpps line rate of a 1 GbE link. Using the Windows or Linux packet layouts finds slightly more services but reduces send rates to 1.389 and 1.276 Mpps, respectively. While filtering packets based on TCP options could be due to defensive mechanisms attempting to block scanning, removing the easily identified static IP ID value of ZMap probes appears to have no impact on scan hit rates. We performed three scans of 10% of IPv4 on TCP/80 in April 2024 with a static IP ID and with a random per-packet IP ID and find that the difference in hit-rate between the random and static IP IDs is not statistically significant. In early 2024, ZMap changed its default behavior to use random per-probe IP IDs.
Figure 7. Hitrates for Varying TCP Options—SYN probes without any TCP options, as originally sent by ZMap, find 1.5–2.0% fewer services on TCP/80 than probes that include options. Mimicking common OSes maximizes coverage. Note truncated $y$ axis.
@@ 5. Lessons and Recommendations
Based on our experiences, we offer several lessons and recommendations for researchers building future Internet measurement tools. These lessons are derived from decisions that we revisited, approached differently in subsequent tools like ZGrab and ZDNS, would make differently if we were to build ZMap today, or believe were fundamental to ZMap’s success. As such, these recommendations are not comprehensive and are inherently opinionated, focusing on decisions where the right choice for ZMap was non-obvious to us at the time. They are, however, the starting point for how we would architect future measurement tools ourselves.
@@
Nearly 50 years later, we cannot agree more with this guidance. It is difficult to predict how researchers will use measurement tools or the environments in which they will operate. ZMap was originally envisioned as a framework where researchers would build customized Scan Modules and Output Modules for service follow up. In practice, the vast majority of researchers use ZMap for service discovery and pipe results to secondary tools for investigation or storage. The output modules ZMap included for connecting to specific databases (e.g., Redis) became liabilities, requiring upkeep and complicating testing and packaging. In time, we removed them, opting to support only Text, CSV, and JSON Lines output.
Recommendation: Build small, simple, easy-to-understand, easy-to-use, and easy-to-test measurement tools that can be creatively assembled, rather than complex applications or software frameworks. Build applications that do one thing well. Continuously output results on a per-record/per-line basis when possible. Avoid proprietary formats and standardize output on well-worn interfaces like CSV, JSON Lines (jsonlines, ), BSON, and Apache Avro. Carefully consider whether binary formats are worth the cost of direct interoperability with existing data processing toolkits and command-line tools.
Usability. ZMap was not the first fast, asynchronous network scanner: Unicornscan (unicornscan, ), Scanrand (scanrand, ), and IRLscanner (leonard2010demystifying, ) were released years prior. IRLscanner was published at IMC; though Unicornscan and Scanrand were both unknown to the team. ZMap was likely more successful than prior tools due to its greater ease of use: it enabled researchers to scan the IPv4 Internet from a single machine by running a single command.
Recommendation: Obsess over ease of installation, usage, and troubleshooting as well as documentation. It is better to have a tool that is easy to use and less full featured than vice versa.
Library and Command Line Wrapper. It is natural to build an application where the command-line interface, application configuration, and operation are intermingled, since most tools are first used via the CLI and this is the path of least resistance. However, this design will limit a tool’s potential to be integrated into larger systems. Automated or continuous measurements, such as the scans that power services like Censys (durumeric2015search, ), are cumbersome to control from the CLI and more suited to a library interface.
@@
Data, Metadata, and Logs. Given the amount of raw data collected by many measurement tools, it is difficult to tell whether experiments are operating as expected without analyzing metadata in real time. In addition, tracking as much information about the execution (e.g., time, version of software, configuration parameters, and environment) helps to later interpret, troubleshoot, or reproduce results. Logs are helpful for human debugging, but they are not inherently designed to be machine-parsable, which is needed for monitoring long-running experiments. Ultimately, we extended ZMap to produce four output streams: (1) data, (2) logs, (3) real-time updates (e.g., packets sent, received, dropped per second), and (4) machine-readable metadata at completion.
@@
Recommendation: Design measurement tools to produce separate streams of data, metadata, and logs. Do not cross these streams, since this complicates downstream processing. Be liberal in what environment and execution information is included in scan metadata, as it is difficult to know a priori what will be useful. Adopt a logging library that supports multiple log levels, and use debug-level logging liberally to enable future troubleshooting. In slight contrast to SoMeta (sommers2017automatic, ), we recommend that metadata collection should be built into measurement tools to maximize ease of use.
@@
Static Types and Output Schema. JSON and CSV provide considerable flexibility for encoding data. For example, JSON objects can have dynamic keys and value types across records. However, downstream applications/databases often do not support this flexibility, and it is easy to create valid but painful to process records.
@@
Recommendation: Define a schema for the data you output. Ensure that each field uses a single, well defined type and that the type of one field does not depend on the value of another field. Avoid maps with dynamic keys, and instead use lists of a static document type. Consider using a tool like JSON Schema (json-schema, ) or ZSchema (zschema, ) to document the structure of output data and metadata.
@@
Versioning and Releases. We released ZMap far too infrequently, repeatedly wanting to include one more feature or bugfix in each release. In particular, ZMap 3.0 was released nearly eight years after the previous release of ZMap 2.1.1. Unfortunately, this meant that most users were either using long out-of-date releases or unversioned code. This made debugging difficult and prevented users from describing the ZMap version they used when publishing.
@@
Recommendation: Follow the Semantic Versioning Specification (semver, ) religiously. Focus on making regular, stable, versioned releases rather than trying to finish a preset amount of work.
@@
Language Choice. When we wrote ZMap, C/C++ were the only practical high-performance systems languages available. It is easy to convince oneself that it is possible to safely write C code; empirical evidence overwhelmingly says the opposite (gaynor2020science, ; whitehouse2024memory, ). Network parsers are particularly hard to implement safely and must protect against attacker-controlled input (chromeruleoftwo, ). ZMap has had multiple regressions that caused incomplete measurements and memory safety bugs, which could have been avoided (e.g., (bug1, ; bug2, ; bug3, )). We also found that memory safety concerns make it harder to review external contributions for safety and correctness, which has reduced the rate at which we merge improvements. If were to implement ZMap today, we would do so in Rust.
@@
Recommendation: Develop tools in modern, memory-safe languages like Rust and Go. While Rust has a relatively steep learning curve, its safety and performance make it ideal for performance-critical applications. Go’s simple syntax and parallelism-oriented architecture make it particularly suited for quickly developing high-performance measurement tools (e.g., ZGrab (zgrab2, ) and ZDNS (zdns, )).
6. Internet Citizenship
Our work provides an opportunity to revisit our original best practices for conducting scans (durumeric2013zmap, ). Overall, we believe that the 2013 recommendations remain a sound set of considerations. However, we encourage the measurement community to treat these recommendations as good practices for most research, not as a set of requirements nor as the basis for having conducted research ethically. For example, there may be situations when scans are better performed unattributed or when opting out specific networks could invalidate results (e.g., tracking a specific threat actor). We additionally recommend that researchers:
@@ 6. Internet Citizenship
    Investigate whether existing datasets (e.g., from Censys or Rapid7) suffice. Often, these datasets provide better coverage than one-off scans and reduce aggregate bandwidth.
@@
    Validate how handshakes will appear in logs. For example, some benign SSH handshakes inadvertently show up as failed authentication attempts that conern operators.
Institutions have adopted different practices for validating opt-out requests (cant-stop-the-scan, ). We have found it necessary to verify the authenticity of exclusion requests. Given that IP address ownership changes over time, it likely makes sense to eventually expire opt-out requests and it may not make sense for institutions to share blocklists. Our team expires requests after 1–2 years and found that in the vast majority of exclusions are not re-requested. We offer the following, updated set of best practices as a recommended starting point when conducting active measurements:
@@
    Minimize Internet Impact. While Internet scanning is a powerful research methodology, it can also affect systems and create work for operators. Consider whether existing open source datasets provide the data you need. If you do perform scans, conduct scans no larger or more frequent than necessary and at the minimum scan rate needed for your research objectives. Publish any scan data you collect.
@@
    Signal Intent. When possible, publish reverse DNS entries, IP WHOIS records, and a website that describes the scans. Ensure that operators can easily contact the research team.
@@
    Provide An Opt-Out Mechanism. Provide a simple mechanism for operators to request exclusion from future scans. Indicate the IP ranges you use for scanning so that operators can drop research traffic themselves.
@@
    Proactively Investigate Effects. Run newly developed scanning code against your own systems to ensure that you understand how scans might affect devices and appear in logs. Start with small experiments before completing full scans in case your scanner causes unexpected problems.
@@
    Coordinate Locally. Coordinate with your local IT and security teams to reduce the risk of overwhelming local networks, as well as to ensure that they know how to handle any inbound inquiries from operators.
@@
    Disclose Results. When appropriate, consider how you can improve the security of the systems you have scanned. Responsibly disclose security problems you uncover and consider notifying vulnerable system owners.
7. Conclusion
The most exciting aspect of building ZMap has been watching how other researchers have used it in unpredicted but valuable ways to meaningfully improve the Internet. We are sincerely thankful to everyone who has contributed, pushing the tool close to feature completion. While the development of new ZMap features has slowed, we are excited to continue to expand the ecosystem of tools that work with ZMap (e.g., ZDNS (zdns, ) and ZGrab (zgrab2, )) and to enable an even broader range of measurement uses. Many of the lessons we learned from maintaining ZMap may seem obvious today, but were not obvious at the time. We hope that our retrospective analysis will help the community build an even richer and more reliable ecosystem of Internet measurement tools moving forward.
Acknowledgements.
We thank all of the contributors to ZMap, the IT and Security teams at the University of Michigan and Stanford University, Michael Bailey, Jeff Cody, and Liz Izhikevich. We thank the reviewers at IMC for their suggestions and feedback. This work was supported by the National Science Foundation under grants CNS-2223360, CNS-1518888, and CNS-2319080, as well as a Sloan Research Fellowship. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of their employers or the sponsors.
Appendix A Ethics
Our work is primarily a metareview of prior studies and a presentation of lessons learned from ZMap. When conducting our own experiments to validate changes made to ZMap, we followed the original guidelines set forth by Durumeric et al. in 2013 (durumeric2013zmap, ). We provide updated recommendations on how to best conduct Internet scanning in Section 6.